OpenStudy (anonymous):

Easy question but need help . . . Of 410 consumers polled, some purchase music on CDs, some only download music, and some do both. If 290 people polled purchase CDs and 220 purchase CDs and download music, how many people download music? A) 340 B) 120 C) 70 D) 220

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

220 Both 290 - 220 = 70 just CD 410 - 220 - 70 = 120 Just download

OpenStudy (danjs):

That is my logic

OpenStudy (danjs):

You could create one of those circle diagrams i think for this,, venn diagram i think its called

OpenStudy (danjs):



OpenStudy (danjs):

Start with 290 total purchased CD, but 220 purchased both, So ONLY CD = 290 - 220 = 70 people If the total is 410 people, then the people that ONLY downloaded is Just DL = total - Both - just CD = 410 - 220 - 70 Just DL = 120

OpenStudy (anonymous):

so the answer is 120 ?